OpenStudy (anonymous):

what are all the real zeros y= (x+7)^3 +2? A) -2- 7 B)-7+2 C)2-7

OpenStudy (anonymous):

(x+7)^3 + 2 = 0 (x+7)^3 = -2 x + 7 = -(2 ^ (1/3)) x = -(2^(1/3)) -7

OpenStudy (anonymous):

:D